OpenStudy (anonymous):

Find the coordinates of the point on the line y = 7x + 3 that are closest to its origin.

OpenStudy (anonymous):

work out x and y-intercepts

OpenStudy (amistre64):

define "closest to its origin" please

OpenStudy (amistre64):

i assume you want a perpendicular line that includes the point (0,0)

OpenStudy (anonymous):

oh yes - i assumed the points on the axes!!

OpenStudy (amistre64):

then solve them in a system of equations that results in the common point

OpenStudy (anonymous):

perpendicular distance = 3/(50)^1/2

OpenStudy (anonymous):

i'm not sure better check it

OpenStudy (amistre64):

we could do vector to maybe :)

OpenStudy (anonymous):

perpendicular satisfying 0,0

OpenStudy (anonymous):

closest to its origin as close as the "x values being a negative decimal not exceeding -3 and the y value can possible range until -17

OpenStudy (amistre64):

so this slope is a "7" ; flip it spank it and throw it back down for the perp... (-1/7)x and that should be it for that ...... but what does this new stuff mean?

OpenStudy (anonymous):

it was a hint at the answer. idk thats what i got

OpenStudy (anonymous):

I was given*

OpenStudy (anonymous):

the closest distance should be \[\frac{3}{\sqrt{50}} \]

OpenStudy (amistre64):

im gonna assume its textbook noise then :) y = 7x + 3 y = (-1/7)x ; multiply by 49 y = 7x + 3 49y = -7x --------------- 50y = 3 y = 3/50 in this case ? maybe?

OpenStudy (anonymous):

-.42, .06

OpenStudy (amistre64):

y = 7x + 3 ; multiply by -1 y = (-1/7)x -y = - 7(7)/7 x -3 y = -1/7 x -------------- 0 = -50/7 x -3 x = -3(7)/50 = -21/50

OpenStudy (amistre64):

:) accursed fraction lol; but ill assume its right

OpenStudy (anonymous):

It is (: As I beat you to the second point :p